# Asymmetrical temporal dynamics of topological edge modes in Su-Schrieffer-Heeger lattice with Kerr nonlinearity

### v1

@@ Abstract
Optical bistability and oscillating phases exist in a Sagnac interferometer and a single ring resonator made of $\chi^{(3)}$ nonlinear medium where the refractive indices are modulated by the light intensity due to the Kerr nonlinearity. An array of coupled nonlinear ring resonators behave similarly but with more complexity due to the presence of the additional couplings. Here, we theoretically demonstrate the bifurcation of topological edge modes which leads to optical bistability in the Su-Schrieffer-Heeger lattice with the Kerr nonlinearity. Additionally, we demonstrate periodic and chaotic switching behaviors in an oscillating phase resulting from the coupling between the topological edge mode and bulk modes with different chiralities, i.e., clockwise and counter-clockwise circulations.
@@ I Introduction
Asymmetrical states emerge when a system lose the balance and thus a symmetry between different components is broken. This can lead to bistability, where the system has two stable states for a single excitation. In some cases, asymmetrical dynamic states can emerge with periodic or chaotic oscillatory behaviors. In photonic systems, optical bistability can appear when the light transmits through a cavity with a nonlinear medium leading to two different optical states, where one mode is dominant (switched on) and the other is quenched (switched off) [1]. For instance, the stable symmetry breaking has been studied for counter-propagating light beams in a Sagnac interferometer [2] and micro-resonator with Kerr nonlinearity [3, 4, 5, 6, 7]. More interestingly, nonlinear optical ring resonators can present rich temporal dynamics with oscillatory behaviors, displaying various types of mode switching, such as chaotic, periodic, and self-switching dynamics [8]. These various dynamics are the result of the nonlinear interaction between the counter-propagating modes and they manifest the symmetry breaking, i.e., unequal intensities of the two counter-propagating modes.
@@
Recently, topologically protected modes have been widely studied in photonic systems due to their intriguing properties, such as unidirectional light propagation and robustness to defects and disorders [9, 10]. In particular, topological edge modes in a one-dimensional (1D) Su-Schrieffer-Heeger (SSH) configuration and two-dimensional (2D) photonic quantum spin-Hall or quantum valley-Hall structures have been employed in an array of coupled lasers, namely topological lasers [11, 12, 13, 14, 15, 16]. Moreover, nonlinear topological photonics has been studied in various platforms such as waveguide arrays [17, 18, 19, 20], microcavity polariton systems [21] and optical resonators [22, 23, 24, 23, 25, 26, 27, 28]. The phase diagrams of a nonlinear SSH model and nonlinear breathing kagome model were drawn for the nonlinear parameters and coupling coefficients between sites [26]. Also, the edge solitons have shown to be stable at any energy when the ratio between the weak and strong couplings falls below a critical value [29]. Up until now, however, no research has demonstrated spontaneous symmetry breaking coming from the nonlinear response for the edge modes in photonic topological insulators.
@@
In this paper, we theoretically show that we can observe asymmetrical temporal dynamics, including optical bistabilities and oscillation phases for topological edge modes in a nonlinear 1D SSH model. The system consists of an array of coupled ring resonators with the Kerr nonlinearity. Using the Lugiato-Lefever equation [30] with additional nearest neighbor couplings, we demonstrate the optical bistability of the topological edge mode in the nonlinear SSH lattice. Finally, we use Poincaré section plots, composed of the maxima of the oscillating intensities, to display the oscillation phases featuring periodic and chaotic switching.
@@ II Linear SSH model with two counter-propagating modes
$\mathbf{H}_{0}=i(\omega_{0}+i\gamma_{0})\mathbb{I}_{M}$ (5)
with $\mathbb{I}_{M}$ the $(M\times M)$ identity matrix. The ring-to-ring coupling is expressed as:
@@ III Lugiato-Lefever equation for nonlinear SSH model
For simplicity, we use the normalized Lugiato-Lefever equation to describe the field amplitudes in our nonlinear SSH model [30]. Notably, the equation is equivalent to the one derived from the temporal coupled-mode theory (see the Appendix A for more details). With the time-varying envelope amplitudes $\tilde{a}(t)$, $\tilde{b}(t)$, defined as $a(t)=\tilde{a}(t)e^{i\omega t}$, $b(t)=\tilde{b}(t)e^{i\omega t}$ respectively, the Lugiato-Lefever equation for a 1D SSH array of nonlinear ring resonators can be written as:
@@
$\mathbf{H}_{0}=-(1+i\eta\Delta)\mathbb{I}_{M},$ (10)
@@
$\mathbf{H}_{NL}=i\eta\mathbb{I}_{M}\times\begin{pmatrix}A|\tilde{a}_{1,+}|^{2}+B|\tilde{a}_{1,-}|^{2}\\ A|\tilde{b}_{1,-}|^{2}+B|\tilde{b}_{1,+}|^{2}\\ A|\tilde{a}_{2,+}|^{2}+B|\tilde{a}_{2,-}|^{2}\\ A|\tilde{b}_{2,-}|^{2}+B|\tilde{b}_{2,+}|^{2}\\ \vdots\\ A|\tilde{a}_{1,-}|^{2}+B|\tilde{a}_{1,+}|^{2}\\ A|\tilde{b}_{1,+}|^{2}+B|\tilde{b}_{1,-}|^{2}\\ A|\tilde{a}_{2,-}|^{2}+B|\tilde{a}_{2,+}|^{2}\\ A|\tilde{b}_{2,+}|^{2}+B|\tilde{b}_{2,-}|^{2}\\ \vdots\end{pmatrix}.$ (11)
To obtain the temporal evolution of the amplitudes for nonlinear SSH lattice, we can solve the time-dependent equation (Eq.(3)) with this Hamiltonian. However, we cannot solve the equation in the frequency domain by simply solviing an eigenvalue equation because it is a system of nonlinear equations.
Figure 2: (a) Optical bistability for seven ring resonators ($M=7$) with detuning $\Delta=1.85$, $v=3$ and $w=7$. (b) Optical bistability with the Kerr nonlinearity in the first ring resonator only. (c), (d) The distributions of intensity for the input intensities corresponding to the dashed vertical lines in (a)
@@ IV Optical bistability of topological edge modes
Optical bistability in a single ring resonator is a result of the XPM between two counter-propagating modes [34]. This means that the Kerr nonlinearity leads to a shift in the resonance frequency of the two counter-propagating modes due to both SPM and XPM, and the coupling via XPM between them leads to spontaneous symmetry breaking above a certain threshold pump intensity [3]. Here, we want to address the question whether we can observe the optical bistability using an edge mode in a nonlinear SSH lattice model.
To theoretically observe the optical bistability in the nonlinear SSH lattice, we consider a SSH array of seven ring resonators ($M=7$) with the nonlinear parameter and the detuning in Ref. [34] and the alternating coupling coefficients ($v=3$ and $w=7$). In our simulations, we scan the pump intensity $I_{s}$ for a certain interval with random initial conditions. Indeed, as shown in Fig. 2(a), we observe the optical bistability in the range between $\log(I_{s}+1)=1.9$ and $\log(I_{s}+1)=2.9$ of pump intensity. Note that the field amplitudes are relatively large for odd sites only (sublattice A), and the intensity decreases exponentially along the right direction for both single stable (Fig. 2(c)) and bistable cases (Fig. 2 (d)), which is the reminiscence of the zero-energy edge modes.
@@
Figure 3: Poincaré section of the maxima of the CW and CCW intensity time series for the first ring resonator in 1D SSH lattice composed seven ring resonators with alternating coupling, $v=3$ and $w=7$. (a) Maximum intensity curves for low input intensities $I_{s}=0.05$ (cyan), $0.1$ (green), $0.2$ (blue), $0.4$ (red), $0.6$ (yellow), $0.8$ (brown). (b) For high input intensities $I_{s}=2$ (red), $6$ (green), $10$ (blue), $14$ (cyan), $16$ (yellow), $20$ (brown), both optical bistability and oscillation regions appear. The vertical dashed lines indicate frequencies determined by the eigenvalue equation for the linear case (Eq.(7)).
@@
To explain the origin of the observed optical bistability, we hypothesize that the optical bistability comes from the symmetry breaking in the first ring only. First, optical bistability in a single ring resonator can occur when the pump intensity is above a certain threshold, called a bifurcation point. This means the first ring will show the optical bistability first as we increase the pump intensity under an excitation close to the zero-energy frequency. Indeed, the detuning $\Delta=1.85$ is smaller than the topological band gap ($82|v-w|=$ 8) meaning the zero-energy edge mode is dominantly excited even though it is off-resonance. This is supported by the field intensity distribution in Fig. 2(d). As the intensities in the rest of rings are much smaller than the first ring (Fig. 2(d)), only the first ring introduces bistability and the modes in the first ring couple to the other rings successively instead of having additional optical bistability from the rest of the rings. Second, to confirm this propagation, we consider the Kerr effect only in the first ring resonator but keep all other parameters the same. This is equivalent to switching off the Kerr effect in the $2N$ ring resonators except the first ring resonator in our original setting. As shown in Fig. 2(b), the intensity-intensity curve has almost identical shape as the original one except slight reduction in the range of $I_{s}$ and slight change in the difference between two counter-propagating mode intensities.
@@
Figure 4: Poincaré sections of the maxima of oscillating coupled intensity as a function of detuning for input intensity $I_{s}=20$, for the first ring resonator from the 1D SSH array. The Poincaré sections for coupling coefficients ($v=3$, $w=7$) and XPM strength ($B$) of 4 and 7 in (a) and (b), respectively. The red shading indicates a symmetric case, the yellow indicates optical bistability, and the cyan indicates oscillations. The vertical dashed lines refer to resonance frequencies for the linear 1D SSH lattice.
@@
Figure 5: Snapshots of intensity distribution in the nonlinear 1D SSH lattice for the CW modes in the left column and the CCW modes in the right column for different values of detuning with $I_{s}=20$, $B=4$.
@@
Figure 6: Time series of intensity and their phase space trajectories for $A=1$, $B=4$ and $I_{s}=20$ at different values of detuning, for the first ring resonators from 1 D SSH array of 7 rings. (a), (b) Optical bistability phase with $\Delta=5.51$. (c), (d) Oscillations without overlapping trajectories. (e), (f) Periodic switching for $\Delta=5.93$. (g), (h) Chaotic switching with $\Delta=7.73$
@@ V Asymmetrical temporal dynamics
Now, let us look at the temporal evolution of the optical intensities of a 1D SSH array that contains seven ring resonators. To visualize oscillation and chaotic phases in our nonlinear system, we will use the Poincaré section obtained by plotting all the local maxima in a time series of oscillating intensities [8]. Since the intensity of each ring in the 1D SSH array follows the same pattern as the intensity of the first ring (see Fig. A1 in Appendix B), we plot the Poincaré sections for the first ring resonator only.
@@
As shown in Fig. 3 (a)(b), the nonlinear SSH lattice exhibits both bistability and oscillation phases in the range of detuning corresponding to the edge mode and two bulk modes with positive detuning for the linear SSH lattice (denoted as the vertical dashed lines). Here, we set $A=1$ and $B=4$ and change the input intensity from 0.05 to 20 denoted with different colors. As we can see in the zoomed view of the plots, these spectra show seven resonance modes; one edge mode with the largest intensity in the middle and six bulk modes on both sides of the edge mode having three on each side. For low input intensities (Fig. 3(a)), the edge mode shifts dramatically and its intensity increases significantly, whereas the bulk modes shift less and their intensities increase slightly. This is due to the localization of intensity at the first ring resonator. For high input intensities (Fig. 3(b)), the CW and CCW modes for edge mode undergo an interaction between them via XPM, leading to an optical bistability. Remarkably, the high input intensity leads to the interaction between the edge mode and the bulk mode near $\Delta=5.2$, resulting in a series of oscillation phases occurring for both CW and CCW modes.
The range of detuning of asymmetrical phases also depends on the XPM strength. Figure 4 (a) and (b) compare the Poincaré sections for two different values of XPM strength $B$. One can observe that increasing $B$ leads to a larger range of detuning for asymmetrical phases including optical bistability (yellow) and oscillation phases (blue). The Poincaré section as a function of XPM strength can be found in Fig.  A2 in Appendix C.
@@
To better understand the asymmetrical dynamic modes, we show the spatial distributions and temporal changes of the excited mode intensity for different detuning in Fig. 5 and Fig. 6. Here, we focus on the case of $I_{s}=20$ and $B=4$ as an example of high input intensity. For the optical bistability ($\Delta=4.77$) shown in Fig. 5(a),(b), both CW and CCW modes have contrasting intensity values, whereas their profiles are similar to the zero-energy edge mode’s profile in a linear SSH model with exponentially decaying non-zero odd-site intensities and zero even-site intensities. The deviations can be attributed to the off-resonance excitation and the interaction between CW and CCW modes via the nonlinear process (XPM). Note that the excited mode is stable as they have constant intensities and appear as two separate points its phase space (Fig. 6 (a),(b)). When we increase the detuning further to $\Delta=5.51$, both CW and CCW mode profile deviates further away from the zero-energy edge mode but the CCW mode profile deviates less still having low intensities at even sites (Fig. 5(c),(d)). Here, the largest value at the first site is related to the zero-energy edge mode and also due to the fact we are exciting the ring resonators from the waveguide on the left side. The dynamics for this detuning (Fig. 6 (c)(d)) is periodically oscillatory, resulting in two distinct regions in the phase space meaning the CW mode intensity is always larger than the CCW mode intensity (the trajectory for CW is further away from the origin). For slightly larger detuning of $\Delta=5.93$, the two trajectories are merged into one meaning that the intensities between the two modes alternates. In the phase space, they are in different two points with the $\pi$ phase difference in the same trajectory. For a large detuning of $\Delta=7.73$, we see chaotic oscillations showing two separate trajectories covering a similar region in the phase space (Fig.  6(g),(h)).
@@
In contrast to the optical bistability coming from the coupling between two counter-propagating modes via nonlinearity, the emergence of the periodic and chaotic oscillations come from the coupling between the edge mode and the bulk modes. The reasoning is below. First, the resonance frequency shift of the edge mode when increasing the intensity is much larger than the ones for the bulk modes as shown in Fig.  4. Second, although the intensity distributions for CW and CCW are close to the bulk modes, there are clear signatures of the edge modes, i.e., an exponentially decaying odd-site intensities and nearly zero even-site intensities (Fig. 5 (d)(f)(g)). Thus, our numerical simulations confirm that the edge mode overlap with bulk mode due to the Kerr nonlinearity results in the periodic and chaotic oscillations.

### v2

# Asymmetrical temporal dynamics of edge modes in Su-Schrieffer-Heeger lattice with Kerr nonlinearity

@@ Abstract
Optical bistability and oscillating phases exist in a Sagnac interferometer and a single ring resonator made of $\chi^{(3)}$ nonlinear medium where the refractive indices are modulated by the light intensity due to the Kerr nonlinearity. An array of coupled nonlinear ring resonators behave similarly but with more complexity due to the presence of the additional couplings. Here, we theoretically demonstrate the bifurcation of edge modes which leads to optical bistability in the Su-Schrieffer-Heeger lattice with the Kerr nonlinearity. Additionally, we demonstrate periodic and chaotic switching behaviors in an oscillating phase resulting from the coupling between the edge mode and bulk modes with different chiralities, i.e., clockwise and counter-clockwise circulations.
@@ I Introduction
Asymmetrical states emerge when a system lose the balance and thus a symmetry between different components is broken. This can lead to bistability, where the system has two stable states for a single excitation. In some cases, asymmetrical dynamic states can emerge with periodic or chaotic oscillatory behaviors. In photonic systems, optical bistability can appear when the light transmits through a cavity with a nonlinear medium leading to two different optical states, where one mode is dominant (switched on) and the other is quenched (switched off) [1]. For instance, the stable symmetry breaking has been studied for counter-propagating light beams in a Sagnac interferometer [2]and micro-resonator with Kerr nonlinearity [3, 4, 5, 6, 7] and also for a singular direction of input light in a 1D chain of ring resonators [8]. More interestingly, nonlinear optical ring resonators can present rich temporal dynamics with oscillatory behaviors, displaying various types of mode switching, such as chaotic, periodic, and self-switching dynamics [9]. These various dynamics are the result of the nonlinear interaction between the counter-propagating modes and they manifest the symmetry breaking, i.e., unequal intensities of the two counter-propagating modes.
@@
Topological phases of matters are classified depending on the symmetry and dimension, featuring topological defects and gapless modes in topological insulators and superconductors [10, 11]. Further, three-dimensional topological band insulator can have dislocation-line modes [12]. More recently, topologically protected modes have been widely studied in photonic systems due to their intriguing properties, such as unidirectional light propagation and robustness to defects and disorders [13, 14]. In particular, topological edge modes in a one-dimensional (1D) Su-Schrieffer-Heeger (SSH) configuration and two-dimensional (2D) photonic quantum spin-Hall or quantum valley-Hall structures have been employed in an array of coupled lasers, namely topological lasers [15, 16, 17, 18, 19, 20]. Moreover, nonlinear topological photonics has been studied in various platforms such as waveguide arrays [21, 22, 23, 24], microcavity polariton systems [25, 26] and optical resonators [27, 28, 29, 28, 30, 31, 32, 33]. The phase diagrams of a nonlinear SSH model and nonlinear breathing kagome model were drawn for the nonlinear parameters and coupling coefficients between sites [31]. Also, the edge solitons have shown to be stable at any energy when the ratio between the weak and strong couplings falls below a critical value [34]. Up until now, however, no research has demonstrated spontaneous symmetry breaking coming from the nonlinear response for the edge modes in photonic topological insulators.
@@
In this paper, we theoretically show that we can observe asymmetrical temporal dynamics, including optical bistabilities and oscillation phases for edge modes in a nonlinear 1D SSH model. The system consists of an array of coupled ring resonators with the Kerr nonlinearity. Using the Lugiato-Lefever equation [35] with additional nearest neighbor couplings, we demonstrate the optical bistability of the topological edge mode in the nonlinear SSH lattice. Finally, we use Poincaré section plots, composed of the maxima of the oscillating intensities, to display the oscillation phases featuring periodic and chaotic switching.
@@ II Linear SSH model with two counter-propagating modes
$\mathbf{H}_{0}=i(\omega_{0}+i\gamma_{0})\mathbb{I}_{2M}$ (5)
with $\mathbb{I}_{2M}$ the $(2M\times 2M)$ identity matrix. The ring-to-ring coupling is expressed as:
@@ III Lugiato-Lefever equation for nonlinear SSH model
For simplicity, we use the normalized Lugiato-Lefever equation to describe the field amplitudes in our nonlinear SSH model [35]. Notably, the equation is equivalent to the one derived from the temporal coupled-mode theory (see Appendix A for more details). With the time-varying envelope amplitudes $\tilde{a}(t)$, $\tilde{b}(t)$, defined as $a(t)=\tilde{a}(t)e^{i\omega t}$, $b(t)=\tilde{b}(t)e^{i\omega t}$ respectively, the Lugiato-Lefever equation for a 1D SSH array of nonlinear ring resonators can be written as:
@@
$\mathbf{H}_{0}=-(1+i\eta\Delta)\mathbb{I}_{2M},$ (10)
@@
$\mathbf{H}_{NL}=i\eta\mathbb{I}_{2M}\times\mathbf{diag}\begin{pmatrix}\mu_{1,a}^{cw},&\mu_{1,b}^{cw},&\ldots,&\mu_{1,a}^{ccw},&\mu_{1,b}^{ccw},&\ldots\end{pmatrix},$ (11)
where
$\mu_{n,a}^{cw}=A|\tilde{a}_{n,+}|^{2}+B|\tilde{a}_{n,-}|^{2},$ (12)
and
$\mu_{n,a}^{ccw}=A|\tilde{a}_{n,-}|^{2}+B|\tilde{a}_{n,+}|^{2}.$ (13)
Similarly we can define $\mu_{n,b}^{cw}$ and $\mu_{n,b}^{ccw}$. Because of the nonlinear terms, the time-dependent equation (Eq.(3)) cannot be changed into an eigenvalue equation. Instead, we can obtain the temporal evolution of the amplitudes $\tilde{a}_{n,\pm}$, $\tilde{b}_{n,\pm}$ by solving it numerically. In this work, we use the Runge-Kutta fourth order method to integrate Eq.(3) with respect to the time.
Figure 2: (a) Optical bistability for seven ring resonators ($M=7$) with detuning $\Delta=2.1$, $A=1$, $B=2.5$, $\gamma_{c}=1$, $v=3$ and $w=7$. (b) Optical bistability with the Kerr nonlinearity in the first ring resonator only. (c), (d) The distributions of intensity for the input intensities corresponding to the dashed vertical lines in (a).
@@ IV Optical bistability of topological edge modes
Figure 3: Poincaré section of the maxima of the CW and CCW intensity time series for the first ring resonator in 1D SSH lattice composed seven ring resonators with alternating coupling, $v=3$ and $w=7$, here $A=1$, $B=4$ and $\gamma_{c}=1$. (a) Maximum intensity curves for low input intensities $I_{s}=0.05$ (yellow), $2$ (cyan), $3$ (green), $4$ (blue), $5$ (red). (b) For high input intensities $I_{s}=20$ (yellow), $25$ (cyan), $30$ (green), $35$ (blue), $40$ (red), both optical bistability and oscillation regions appear. The vertical dashed lines indicate frequencies determined by the eigenvalue equation for the linear case (Eq.(7)).
Optical bistability in a single ring resonator is a result of the XPM between two counter-propagating modes [39]. This means that the Kerr nonlinearity leads to a shift in the resonance frequency of the two counter-propagating modes due to both SPM and XPM, and the coupling via XPM between them leads to spontaneous symmetry breaking above a certain threshold pump intensity [3]. Here, we want to answer the question whether we can observe the optical bistability using an edge mode in a nonlinear SSH lattice model.
To theoretically observe the optical bistability in the nonlinear SSH lattice, we consider a SSH array of seven ring resonators ($M=7$) with detuning of $2.1$, $A=1$, $B=2.5$ and $\gamma_{c}=1$ as and the alternating coupling coefficients ($v=3$ and $w=7$). In our simulations, we vary the pump intensity $I_{s}$ for a certain interval with random initial conditions. Indeed, as shown in Fig. 2(a), we observe the optical bistability where the symmetry between CW and CCW modes is broken for the pump intensity range between $\log(I_{s}+1)=3.4$ and $\log(I_{s}+1)=5.9$. This means that one of the CW and CCW modes becomes stronger while the other mode is suppressed. As shown in Fig. 2 (c, d), the amplitude decreases gradually towards the bulk of the lattice system. Note that the amplitude is relatively large for odd sites only (sublattice A), and the intensity decreases exponentially along the right direction for both single stable (Fig. 2(c)) and bistable cases (Fig. 2 (d)), which is the reminiscence of the zero-energy edge modes.
@@
To explain the origin of the observed optical bistability, we hypothesize that the optical bistability comes from the symmetry breaking in the first ring only. First, optical bistability in a single ring resonator can occur when the pump intensity is above a certain threshold called a bifurcation point. This means the first ring will show the optical bistability first as we increase the pump intensity under an excitation close to the zero-energy frequency. Indeed, the detuning $\Delta=1.85$ is smaller than the topological band gap ($82|v-w|=$ 8) meaning the zero-energy edge mode is dominantly excited even though the system is at off-resonance. This is supported by the field intensity distribution in Fig. 2(d). As the intensities in the rest of rings are much smaller than the first ring (Fig. 2(d)), only the first ring introduces bistability and the modes in the first ring couple to the other rings successively instead of having additional optical bistability from the rest of the rings. Second, to confirm this propagation of asymmetric intensities, we consider the Kerr effect only in the first ring resonator but keep all other parameters the same. This is equivalent to switching off the Kerr effect in the $2N$ ring resonators except the first ring resonator in our original setting. As shown in Fig. 2(b), the intensity-intensity curve has almost identical shape as the original one except slight reduction in the range of $I_{s}$ and slight change in the difference between two counter-propagating mode intensities.
@@
Figure 4: (a) Poincaré sections of the maxima of oscillating coupled intensity as a function of detuning for pump intensity $I_{s}=40$, for the first ring resonator from the 1D SSH array. The Poincaré sections for coupling coefficients ($v=3$, $w=7$) and XPM strength ($B$) of 4. The red shading indicates a symmetric case, the yellow indicates optical bistability, and the cyan indicates oscillations. The vertical dashed lines refer to resonance frequencies for the linear 1D SSH lattice. (b) Dissymmetry for CW and CCW intensity modes. (c) polarization in A and B sites. (d) Participation ratio for intensity.
@@
Figure 5: Snapshots of intensity distribution in the nonlinear 1D SSH lattice with $I_{s}=40$, $B=4$, two figures in each row have the same detuning values for the CW modes and the CCW modes, respectively. (a), (b) in the optical bistability regime, (c), (d) in the oscillation regime, and (e), (f) in the oscillation switching regime.
@@
Figure 6: Time series of intensity and their phase space trajectories for $A=1$, $B=4$ and $I_{s}=40$ at different values of detuning, for the first ring resonators from 1 D SSH array of 7 rings. (a), (b) Optical bistability phase with $\Delta=4.77$. (c), (d) Oscillations without overlapping trajectories. (e), (f) Periodic switching for $\Delta=7.51$. (g), (h) Chaotic switching with $\Delta=8.78$.
@@ V Asymmetrical temporal dynamics
Now, let us look at the temporal evolution of the optical intensities of a 1D SSH array that contains seven ring resonators. To visualize oscillation and chaotic phases in our nonlinear system, we will use the Poincaré section obtained by plotting all the local maxima in a time series of oscillating intensities [9]. Since the intensity of each ring in the 1D SSH array follows the same pattern as the intensity of the first ring (see Fig. A1 in Appendix C), we plot the Poincaré sections for the first ring resonator only.
@@
As shown in Fig. 3 (a)(b), the nonlinear SSH lattice exhibits both bistability and oscillation phases in the range of detuning corresponding to the edge mode and two bulk modes with positive detuning for the linear SSH lattice (denoted as the vertical dashed lines). Here, we set $A=1$ and $B=4$ and change the pump intensity from 0.05 to 20 denoted with different colors. As we can see in the zoomed view of the plots, these spectra show seven resonance modes; one edge mode with the largest intensity in the middle and six bulk modes on both sides of the edge mode having three on each side. For low pump intensities (Fig. 3(a)), the edge mode shifts to larger detuning values dramatically and its intensity increases significantly, whereas the bulk modes shift less and their intensities increase slightly. This is due to the localization of intensity at the first ring resonator. For high pump intensities (Fig. 3(b)), the CW and CCW modes for edge mode undergo an interaction between them via XPM, leading to an optical bistability. Remarkably, the high pump intensity leads to the interaction between the edge mode and the bulk mode near $\Delta=6.8$ for $I_{s}=40$ as shown in Fig. 4(a), resulting in a series of oscillation phases occurring for both CW and CCW modes. Note that CW and CCW intensities for bistable and oscillation phases can be exchanged any time because the chirality depends on the history of the dynamic system which is determined by the initial conditions in our case.
To quantify the degree of the CW-CCW symmetry breaking (see Appendix B), we introduce the dissymmetry parameter defined in a similar manner for a single microcavity in Ref. [40] as:
$D=\frac{\sum_{n}I_{CW,n}-\sum_{n}I_{CCW,n}}{\sum_{n}I_{CW,n}+\sum_{n}I_{CCW,n}}.$ (14)
Then, a positive (negative) value of $D$ means the CW (CCW) mode is stronger than the other, corresponding to the optical bistability, and $D=0$ means the mode is CW-CCW symmetric. For the bistability regime, the nonlinear system shows relatively large dissymmetry close to 1 as shown in Fig. 4(b), meaning the CW-CCW symmetry is strongly broken. However, for the oscillations regime ($\Delta>6.9$), the dissymmetry has a random value between -1 and 1 and becomes very sensitive to the detuning. A similar trend can be observed in the dissymmetry calculated for the first ring only, reconfirming that the origin of the optical bistability is mainly due to the bifurcation of CW and CCW modes in the first ring.
To investigate the sublattice symmetry breaking (see Appendix B), we calculate polarization ($P$) in terms of $\sum_{n}I_{A,n}$ (for sublattice A) and $\sum_{n}I_{B,n}$ (for sublattice B) for each direction using the following equation:
$P=\frac{\sum_{n}I_{A,n}-\sum_{n}I_{B,n}}{\sum_{n}I_{A,n}+\sum_{n}I_{B,n}}.$ (15)
Positive polarization values in Fig. 4(c) indicate that the modes for the bistability regime are localized at sublattice A; a value close to 1 confirms that the energy of the modes is concentrated at A sites. This is a reminiscence of topological zero-energy mode which would show a completely polarized distribution, i.e., $P=1$.
To characterize the localization of the intensity of dynamic modes in the whole SSH lattice, we calculate the participation ratio ($PR$), which is given by [41]:
$PR=\frac{(\sum_{n}I_{n})^{2}}{\sum_{n}|I_{n}|^{2}}.$ (16)
A large value of $PR$ refers to delocalization. At small values of detuning($\Delta\leq 1.5$), as shown in Fig. 4(d), the nonlinear SSH lattice exhibits almost identical delocalization behavior for both CW and CCW modes. In contrast, for larger detuning ($\Delta\geq 1.5$), one of CW and CCW modes is more localized that the other.
@@
To better understand the asymmetrical dynamic modes, we show the spatial distributions and temporal changes of the excited mode intensity for different detuning in Fig. 5 and Fig. 6. Here, we focus on the case of $I_{s}=40$ and $B=4$ as an example of high pump intensity. For the optical bistability ($\Delta=4.77$) shown in Fig. 5(a),(b), both CW and CCW modes have contrasting intensity values, whereas their profiles are similar to the zero-energy edge mode’s profile in a linear SSH model with exponentially decaying non-zero odd-site intensities and zero even-site intensities. The deviations can be attributed to the off-resonance excitation and the interaction between CW and CCW modes via the nonlinear process (XPM). Note that the excited mode is stable as they have constant intensities and appear as two separate points in the phase space (Fig. 6 (a),(b)). When we increase the detuning further to $\Delta=7.3$, both CW and CCW mode profile deviates further away from the zero-energy edge mode but the CCW mode profile deviates less still having low intensities at even sites (Fig. 5(c),(d)). Here, the largest value at the first site is related to the zero-energy edge mode and also due to the fact we are exciting the ring resonators from the waveguide on the left side. The dynamics for this detuning (Fig. 6 (c)(d)) is periodically oscillatory, resulting in two distinct regions in the phase space meaning the CW mode intensity is always larger than the CCW mode intensity (the trajectory for CW is further away from the origin). For slightly larger detuning of $\Delta=7.51$ (Fig. 6 (e)(f)), the two trajectories are merged into one meaning that the intensities between the two modes alternate. In the phase space, they are located at different two points with the $\pi$ phase difference in the same trajectory. For a large detuning of $\Delta=8.78$, we see chaotic oscillations showing two separate trajectories covering a similar region in the phase space (Fig.  6(g),(h)).
@@
In contrast to the optical bistability coming from the coupling between two counter-propagating modes via nonlinearities (XPM and SPM), the emergence of the periodic and chaotic oscillations has its origin in the coupling between the edge mode and the bulk modes. The reasoning is as follows: Firstly, the resonance frequency shift of the edge mode when increasing the intensity is much larger than the ones for the bulk modes (Fig.  4(a)), and there are clear signatures of the edge modes, i.e., an exponentially decaying odd-site intensities and nearly zero even-site intensities (Fig. 5 (c-h) although the intensity distributions are getting close to the bulk modes. Thus, our numerical simulations confirm that the edge mode overlap with bulk mode due to the Kerr nonlinearity results in the periodic and chaotic oscillations.